OpenStudy (anonymous):

I NEED HELP ASAP! Medal would be given! C:

OpenStudy (anonymous):

@abb0t @timo86m @AgentZero help guys!

OpenStudy (anonymous):

i have this so far: N=d +5 as an equation but idk what to do or if im right :C

OpenStudy (anonymous):

I know its 7 dimes and 12 nickels but idk how though?

OpenStudy (anonymous):

7x(.10) + 12x(.05) = $1.30 i know this is what im suppose to do but where did the 7 and 12 come from?

OpenStudy (anonymous):

let me dbl check mines :D

OpenStudy (anonymous):

did you delete one of your posts?

OpenStudy (anonymous):

Naw why? 0_o & i get that but how do you ended up with 7 & 12 like how though i know all the plug in stuff and eqution i just dont get why its 7 & 12 besides thats it shows it has 5 more nickels

OpenStudy (anonymous):

5 more nickels than dimes n=.05 d=.1 the 2 above by definition x =5+ y ^ 5 more and we assume y is number of nickels xn+yd=1.3

OpenStudy (anonymous):

I got it now :D

OpenStudy (anonymous):

First step is to clean it up n=.05 eq1 d=.1 eq 2 x =5+ y eq 3 xn+yd=1.3 eq 4 --------------- we plug in n and d to the 4th equation x*.05+y*.1=1.3 eq 5 then we plug in eq 3 into eq 5 (5+y)*.05+y*.1=1.3 .25+.05y + .1y =1.3 distributed the .05 then rearange y*.1 to .1y .25+.15y=1.3 add like terms the ys in this case .15y=1.05 subtract .25 from both sides y= 7 divide both sides by .15 now you have y=7 one of the answers use eq 3 x =5+ y x=5+7 and x = 12